OpenStudy (anonymous):

What is the area of a circle with a radius of 2 meters? Use twenty-two sevenths for pi.

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Area of the circle is pi times the square of the radius...

OpenStudy (anonymous):

Except this time you use 22/7 for pi because the question wants you to...@iceicebaby

OpenStudy (anonymous):

so it would be 12 and 4 over 7 right?

OpenStudy (anonymous):

\[\bf A=2^2 \times \frac{22}{7}=\frac{ 4 \times 22}{7}=\frac{88}{7}=12 \frac{ 4 }{ 7 }\]So that's right... @iceicebaby